How many moles of aluminum are required to completely react with 107 mL of 6.00 M H2SO4 according to the balanced chemical reaction: 

2 Al(s) + 3 H2SO4(aq) → Al2(SO4)3(aq) + 3 H2(g)


How many moles of AgI will be formed when 75.0 ml of 0.300 M AgNO3 completely reacted according to the balanced chemical reaction: 

2 AgNO3(aq) + Cal2(aq) + 2 Agl(s) + Ca(NO3)2(aq)
Answer #2
Answer #1
Answer #3

Ans. Mole of AgNO​​​3 = molarity × volume (L)

Mole of AgNO​​​3 = 0.300 × 75/1000 = 2.25 × 10-2 mol

According to balanced reaction

2 Mole of AgNO​​​3 produced Agl = 2 mole

2.25 × 10-2 mol AgNO​​​3 produced Agl = (2/2) × 2.25 × 10-2 = 0.0225 mol